Factor complete:
a. x^2-3x-10
b. 4z^4 + 12z^3 - 40z^2

Answer by funmath(2933)

Factor complete:
a. x^2-3x-10
(x__)(x__)
The blanks need to have integers that multiply to give you -10, and add to give you -3. -5*2=-10 and -5+2=-3
(x-5)(x+2)
b. 4z^4 + 12z^3 - 40z^2 Factor out the GCF 4z^2
4z^2(z^2+3z-10)
4z^2(z__)(z__)
The blanks need to be filled with numbers that multiply to give you -10, but add to give you 3. 5*-2=-10 5+-2=3
4z^2(z+5)(z-2)
:
Happy Calculating!!!